### v1

# Beyond the Silence: How Men Navigate Infertility Through Digital Communities and Data Sharing

Abstract.
Men experiencing infertility face unique challenges navigating Traditional Masculinity Ideologies that discourage emotional expression and help-seeking. This study examines how Reddit’s r/maleinfertility community helps overcome these barriers through digital support networks. Using topic modeling (115 topics), network analysis (11 micro-communities), and time-lagged regression on 11,095 posts and 79,503 comments from 8,644 users, we found the community functions as a hybrid space: informal diagnostic hub, therapeutic commons, and governed institution. Medical advice dominates discourse (63.3%), while emotional support (7.4%) and moderation (29.2%) create essential infrastructure. Sustained engagement correlates with actionable guidance and affiliation language, not emotional processing. Network analysis revealed structurally cohesive but topically diverse clusters without echo chamber characteristics. Cross-posters (20% of users) who bridge r/maleinfertility and the gender-mixed r/infertility community serve as navigators and mentors, transferring knowledge between spaces. These findings inform trauma-informed design for stigmatized health communities, highlighting role-aware systems and navigation support.
Male infertility, Traditional masculinity ideologies, Online health communities, Digital peer support, Masculinity and reproductive health Emotional coping and stigma, Health-related HCI design, Trauma-informed Design, Infertility discourse on Reddit
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
Male infertility accounts for 20–30% of infertility cases (Agarwal et al., 2015), yet remains under-discussed (Patel et al., 2019; Arya and Dibb, 2016), under-supported (Costa Figueiredo et al., 2017; Patel et al., 2019), and heavily stigmatized (Gannon et al., 2004; Dooley et al., 2011). Beyond thwarting reproductive goals, it challenges cultural linkages between masculinity and virility, introducing biological, relational, and psychological uncertainties while offering few social scripts or institutional resources for coping. Unlike many areas of reproductive health that benefit from gendered solidarity and advocacy, men experiencing infertility often face profound isolation. In the absence of formal or culturally accessible support systems, many turn to online spaces for guidance and connection. This study focuses on one such forum, r/maleinfertility described as a space “for men experiencing infertility and male perspectives on infertility. Partners are encouraged to participate, but we ask they post in the daily Partner’s Perspectives thread.”
Research in HCI and CSCW shows how online health and parenting communities foster emotional expression, identity work, and peer support (Ammari et al., 2018; Andalibi and Forte, 2018; Andalibi, 2020). Studies of stay-at-home fathers and non-normative caregivers reveal that men use online forums to reframe vulnerability and caregiving as compatible with strength and responsibility (Ammari et al., 2017, 2019; Lupton et al., 2016), sometimes creating father-only spaces (Ammari and Schoenebeck, 2016). Yet—with the exception of Patel et al. (Patel et al., 2019)—male infertility has received little attention in social computing scholarship, despite its potential to illuminate gendered stigma, identity repair, and support-seeking. Building on this literature, we examine Reddit as a platform where men disclose and process infertility experiences, asking:
RQ1a: What kinds of discourse emerge in male infertility communities on Reddit?
While female infertility discourse is well theorized in terms of emotional labor and peer support (Patel et al., 2019), male infertility discourse remains underexplored. Understanding its emotional and linguistic structures can inform platform design and peer support. Prior work suggests that affective and narrative strategies—such as vulnerability or mentoring—boost engagement (Ammari et al., 2025; Andalibi, 2020), but it is unclear which forms of self-expression are most accepted or rewarded in male infertility communities. We therefore also ask:
RQ1b: What types of discourse and user attributes lead to higher engagement in these communities?
Online spaces like Reddit operate across macro (platform-wide), meso (multi-subreddit), and micro (single-subreddit or sub-community) levels of norms (Chandrasekharan et al., 2018; Cai and Wohn, 2021; Ammari et al., 2022; Vachher et al., 2020). These norms shape moderation and acceptable discourse, introducing differences even within one subreddit (Weld et al., 2024). Such dynamics matter in stigmatized contexts, where users may encounter narrow perspectives such as despair narratives or medical fatalism (Randazzo and Ammari, 2025, 2023). To investigate these dynamics, we ask:
RQ2a: Are there distinct sub-communities within male infertility subreddit?
RQ2b: How do different sub-communities frame infertility, masculinity, and coping strategies?
RQ2c: Do these sub-communities function as echo chambers?
Emerging HCI research highlights the role of cross-posters—users who bridge multiple communities—as knowledge brokers and empathy-builders (Ammari et al., 2025). In foster care and trauma recovery spaces, cross-posters act as informal mentors, attract more responses, and introduce diverse viewpoints. In male infertility contexts, cross-posting between male-centric and broader infertility communities (e.g., r/infertility) may broaden resources and normalize emotional expression. Yet there remains more to be known about these dynamics:
RQ3: What are the characteristics and functions of users who cross-post between male infertility subreddits and broader infertility communities?
By addressing these questions, we extend Patel et al. (Patel et al., 2019) through large-scale analysis of male infertility communities and contribute to HCI conversations on trauma-informed design, digital support infrastructures, and gendered online identity work. Our findings aim to inform the design of more inclusive, empathetic platforms for reproductive health support, particularly for conditions that remain invisible or culturally silenced.
2. Related Work
In this section, we locate our study at the intersection of gender, online support, and design. We first review how Traditional Masculinity Ideologies (TMI) shape men’s coping and help-seeking, highlighting how infertility constrains identity and disclosure. We then examine how online communities foster connection and identity (re)construction through personal data sharing and peer interaction. Finally, we draw on trauma-informed design to consider platform affordances—pseudonymity, algorithmic curation, and moderation—that enable or hinder safety, agency, and peer support. Together, these strands explain why men turn to Reddit, how support practices form, and which design choices sustain engagement in male-infertility communities.
2.1. The Role of Traditional Masculinity Ideologies in Men’s Mental Health
Men often show distinct behavioral patterns, including higher rates of aggression (Björkqvist, 2018), substance misuse (McHugh et al., 2018), and suicide (Coleman et al., 2020). These patterns are partly rooted in Traditional Masculinity Ideologies (TMI), which valorize self-reliance, emotional restraint, and dominance (Messerschmidt, 2019). By discouraging emotional expression, TMI fosters harmful coping strategies like alcohol use and unmanaged stress (Cleary, 2012; Oliffe et al., 2023; Tsan et al., 2011), which can escalate into aggression, social withdrawal, and suicidal behavior (Krysinska and Lester, 2010).
Men’s emotional expression is context-dependent. Limited social settings like funerals or sporting events can permit displays of emotion without threatening masculine norms (Walton et al., 2004), yet pervasive cultural messages such as “boys don’t cry” continue to suppress emotional development (Frosh et al., 2003; Creighton et al., 2013). Dolan et al. (Dolan et al., 2017) found that men facing infertility often avoided sharing their struggles with friends or partners to escape ridicule, adopting a perverse passivity in which infertility was something endured rather than addressed. Our study explores whether digital spaces—offering anonymity and asynchronous interaction—can create alternative contexts that legitimize emotional expression.
Gender differences are also evident in help-seeking. Women are significantly more likely than men to engage with online mental health therapy (Titov et al., 2017), and men who begin therapy frequently drop out due to stigma, vulnerability, or concerns about autonomy—barriers reinforced by TMI (Englar-Carlson, 2006; Agüera et al., 2017). These obstacles compound psychological distress and encourage maladaptive coping (Seidler et al., 2016).
Within infertility care, help-seeking is further constrained by the framing of infertility as a female-centered issue, with male discussions focusing narrowly on sperm quality and count (Gannon et al., 2004). Gannon observed that declining sperm counts were depicted as a “war” against men’s bodies, reinforcing narratives of emasculation or quasi-castration. As Patel et al. note (Patel et al., 2019, P.10), such portrayals can deter men from seeking help—whether consulting doctors, confiding in friends, or even talking with partners. If virility and reproduction are central TMI ideals, infertility relegates men, even those otherwise “normal,” to the margins of masculinity.
2.2. Digital Pathways to Social Connection, Belonging, and Well-being in Men’s Mental Health
Social connectedness is critical to physical and mental health; its absence raises risks of cardiovascular disease and early mortality (Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022). Men may be especially vulnerable because their social networks are often smaller and less emotionally expressive (Steptoe et al., 2013). In this context, digital platforms have emerged as key venues for connection and support in diverse areas including fatherhood (Ammari et al., 2019, 2017), pregnancy loss (Andalibi, 2019), parenting children with special needs (Ammari and Schoenebeck, 2015), and foster care (Lee et al., 2025; Ammari et al., 2025; Lee et al., 2021; Fowler et al., 2022). We focus on two central functions of these communities: (1) building collective knowledge and informal expertise, and (2) (re)building identity and belonging.
2.2.1. Personal Data Sharing as Collective Knowledge and Informal Expertise
Online support communities allow people to connect with others who share similar experiences (Fox, 2014; Rains and Wright, [n. d.]; Thoits, 2011), offering critical insights and anecdotal evidence to peers and caregivers (Schorch et al., 2016; Soc, 2013). Individuals weigh benefits and drawbacks of data sharing (Wu et al., 2023), but intrinsic rewards such as self-worth and satisfaction motivate ongoing participation (Yan et al., 2016; Soc, 2011).
Self-tracking is common in managing chronic conditions (Li et al., 2010), yet sharing such data can be emotionally fraught (Knittel et al., 2021). Posting personal metrics online provides emotional support, celebrates progress, solicits information, and fosters accountability (Chung et al., 2017; Gui et al., 2017; Lu et al., 2021). These shared narratives help users reinterpret their data, validate experiences, and reinforce coping mechanisms (Knittel et al., 2021). Patel et al. similarly found men sharing detailed fertility regimens and digital narratives of their infertility journeys (Patel et al., 2019, P.9).
For those navigating infertility, chronic illness, or other stigmatized conditions, sharing is more than disclosure: it invites empathy and solidarity (Levonian et al., 2021), normalizes experience, and alleviates shame and uncertainty (Mamykina et al., 2015; Choe et al., 2017). This communal exchange supports emotional healing, motivates continued self-tracking, and co-constructs meaning within online health communities (Jing Wen Lin a, 2023).
2.2.2. (re)Building Identity and Belonging to a Community
A sense of belonging protects against depression and suicidal ideation (McLaren and Challis, 2009), yet strategies for fostering it among socially isolated men are limited. Digital forums connect individuals through shared experience and mutual support (Biagianti et al., 2018). Subjective well-being—spanning emotional stability, social engagement, and personal growth—is similarly vital (Leite et al., 2019), and digital interventions like two-way SMS therapy can rival face-to-face treatment for depression and anxiety (Hull et al., 2020). Effective message design and timing further enhance such interventions (Rahman and Nayan, 2022).
Online health communities extend these benefits by enabling patients and caregivers to provide mutual support (Levonian et al., 2021; Joy et al., 2025), helping men—who often have limited offline networks—combat isolation, find peers, and strengthen relationships (Rains, 2024). For marginalized groups, such communities are essential spaces to organize and receive support, including for stigmatized conditions like obesity (Payne et al., 2025) or mental wellness (Progga et al., 2023). Patel et al. document men explicitly requesting support in these settings (Patel et al., 2019, pp.8-10).
These forums also foster identity reconstruction and psychological resilience, offering safe environments for trauma recovery and emotional validation (Ammari et al., 2025; Randazzo and Ammari, 2025). Designing such spaces requires trauma-informed principles (Randazzo et al., 2023), which emphasize respecting user boundaries and language, avoiding forced disclosure or stigma reproduction, and supporting agency through safe expressive modes like digital storytelling (A. Ankrah et al., 2022). We next discuss key tenets of trauma-informed design.
2.3. Trauma-informed Design
Trauma-informed design frameworks (Chen et al., 2022; Scott et al., 2023) highlight how social media affordances—such as algorithmic recommendation and comment sorting (Parchoma, 2014)—and governance practices like moderation (Jhaver et al., 2018) can support vulnerable populations. Three core principles guide this work: peer support, user choice, and safety, all underpinned by transparency and trust (Bellini, 2023; Dym and Fiesler, 2020).
2.3.1. Peer support: Building supportive online spaces
Trauma-informed design centers shared understanding and mutual aid (Chen et al., 2022; Scott et al., 2023). Peer support is shaped by feedback mechanisms: direct signals like upvotes and replies communicate community norms (LaViolette and Hogan, 2019), while observing others’ posts also teaches these norms (Randazzo and Ammari, 2023). Such signals—positive or negative—serve as barometers of community endorsement.
Algorithmic recommendation can enhance transportability, exposing users to narratives that resonate with their experiences (Randazzo and Ammari, 2023; Green, 2021). Yet algorithmic opacity poses risks: emotional content may be amplified to drive engagement, harming well-being and limiting community control (Matias, 2019; Jhaver et al., 2018; Fabbri, 2023).
2.3.2. Pseudonymity, Algorithmic Curation, and Choice
Reddit’s pseudonymity and limited identity cues facilitate disclosure of stigmatized experiences such as sexual assault, trauma, family struggles, and eating disorders (Andalibi, 2019; Andalibi et al., 2018; Haimson et al., 2015; Randazzo and Ammari, 2023; Chancellor et al., 2018). Users exercise agency by selecting communities (Hwang and Foote, 2021), cross-posting content (Datta et al., 2017), and managing overlapping networks of support. Platform features like badges, awards, and peer recognition further influence participation: awards encourage newcomers, while badges help sustain engagement among committed members (Burtch et al., 2022; Trujillo, 2022; Bornfeld and Rafaeli, 2017).
2.3.3. Moderation and community governance
Safety depends on effective moderation that defines acceptable discourse, removes harmful content, and bans disruptive users (Jhaver et al., 2018; Mansour and Francke, 2021). Automated tools assist at scale (Gillespie, 2018; Jhaver et al., 2019) but often miss contextual nuance (Ammari et al., 2022). Because definitions of harm are context-specific (Gillespie, 2018, p.98), automated systems struggle with “sticky,” emotionally complex content. Algorithmic opacity can further erode trust when moderation decisions remain unexplained (Kuo et al., 2023).
3. Methods
This study is part of a broader project on men’s traumatic experiences online. We began with a two-month survey of eleven infertility-, parenting-, adoption-, and related subreddits to map discursive patterns and identify communities of interest. All four team members read posts, took notes, and exchanged memos.
Focusing on male infertility, we collected all posts and comments from r/maleinfertility via the PushShift API (Baumgartner et al., 2020) and the Arctic Shift dataset (Heitmann, 2025), yielding 79,503 comments and 11,095 posts by 8,644 users. Because roughly 20% of these users (1,725) also post to r/infertility, we analyzed both community dynamics in r/maleinfertility and characteristics of these cross-posting users.
Our approach follows Tornberg and Tornberg’s mixed-methods framework (Törnberg and Törnberg, 2016), combining computational modeling with qualitative interpretation. Qualitative analysis guided topic selection, model refinement, and interpretation.
3.1. Model Training and Parameter Tuning
We applied BERTopic (Grootendorst, 2022) to Reddit threads treated as documents, using BERT embeddings to cluster semantically similar content (Reimers and Gurevych, 2019; Angelov, 2020). To optimize performance we trained 32 models, selecting the best via u_mass coherence (Newman et al., 2010; Röder et al., 2015; Chang et al., 2009). The top configuration (coherence = –0.79) produced 115 topics. For each comment we assigned the dominant topic based on BERTopic’s probability distributions (McInnes et al., 2017; Campello et al., 2013). Full hyperparameter details appear in Appendix A.1.
3.1.1. Qualitative coding
We sampled 15 comments per top topic (including parent threads). Each topic was double-coded by two authors, with regular meetings to define topic boundaries and select exemplar quotes, following consensus methods (McDonald et al., 2019). Through axial coding (Wall and Gannon-Leary, 1999) we identified 14 cross-topic themes—e.g., Advice on Insurance, Doctors, and Medication Expense; Comfort Posting; Partner Emotional Support—which we grouped into three meta-themes: Medical Advice, Emotional Support, and Moderation (see Appendix A.2).
3.2. Time-lagged Regression: Future Use Predictors
We analyzed comment-level records from r/maleinfertility, containing a user identifier and a timestamp. Events were binned into monthly periods. We operationalized community activity as the number of comments per user per month. To account for temporal dynamics, we created lagged predictors that reflected each feature’s value in the prior month. The dependent variable was defined as the number of comments contributed in the subsequent month ($t+1$). Only users with at least two months of data were included.
In earlier CHI and CSCW work, time-lagged regression was used to uncover temporal dynamics in digital behavior. Mitra et al.(Mitra et al., 2017) applied it to model how employee engagement spreads across a corporate social network, using prior engagement levels of an individual’s contacts as lagged predictors of their future engagement. Li et al. (Li et al., 2022) used time-lagged regression to predict the emergence of social media gatekeepers on Twitter by linking users’ past behaviors and profile features to their future influence in immigration-related news dissemination. In both cases, time-lagged models enabled the researchers to infer directional and predictive relationships over time.
3.2.1. Feature Set Permutations
We compared models using user metadata, LIWC categories, and thematic weights (§3.1.1). The best specification (RMSE = 2.88, $R^{2}$ = 0.66) combined all three.
User metadata (11 variables)
: average comment length and sentiment (VADER (Elbagir and Yang, 2019; Hutto and Gilbert, 2014)); number of responses and karma per comment (LaViolette and Hogan, 2019); response probability; whether a comment was edited; average response length, sentiment, and karma; time to first response; and total responses received.
LIWC (70 features)
: lexical and syntactic categories validated for social media (Pennebaker et al., 2001, 2007; De Choudhury et al., 2013).
Themes (14 features)
: mean topic weights aggregated into the qualitative themes (see §4.1).
We standardized all predictors and used L1 (Lasso), L2 (Ridge), and Elastic Net regularization to reduce dimensionality and multicollinearity.
3.3. Studying interactions within the r/maleinfertility community:Network Analysis
We built a directed, weighted reply network where nodes are users and edges represent replies. Edges were aggregated by interaction count, self-loops removed, resulting in 8,536 nodes and 53,081 edges.
3.3.1. Community detection
We symmetrized edge weights and applied the Louvain algorithm to maximize modularity (Newman, 2018; Cohen-Addad et al., 2020). We refer to each of the micro-communities in this dataset as clusters in our analysis to reduce any confusion with the r/maleinfertility community.
3.3.2. Cluster Structural and Linguistic Analysis
Conductance (structural insulation)
The conductance $C$ of a cluster is defined by the total volume of edges leaving it as a ratio of the total edges:
$\phi(C)\;=\;\frac{W_{\text{out}}(C)}{\,W_{\text{in}}(C)+W_{\text{out}}(C)\,}\;\in[0,1].$
Lower $\phi$ implies stronger internal cohesion.
Reciprocity in Networks
captures the proportion of mutual ties within a cluster (Morini et al., 2021)
$r=\frac{\sum_{i\neq j}A_{ij}A_{ji}}{\sum_{i\neq j}A_{ij}}.$
Purity
is the fraction of nodes sharing the most common stance label (theme × sentiment) (Morini et al., 2021; ALDayel and Magdy, 2021; Naskar et al., 2016)
$P_{c}=\prod_{a\in A}\max\left(\sum_{v\in c_{a}}(v)\right)$
Theme entropy
quantified within-cluster topical diversity using Shannon entropy:
$H(C)\;=\;-\sum_{t}p(t\mid C)\,\log p(t\mid C).$
Lower $H(C)$ implies a narrow topical focus (content homogeneity), while higher $H(C)$ indicates more diverse subject matter.
3.3.3. Detecting Echo Chambers
Following Morini et al. (Morini et al., 2021), we define echo chambers as clusters with Conductance ¡ 0.5 and Purity ¿ 0.7. To define purity, we start by labeling user stance which we defined as (Qualitative Theme — VADER sentiment). In other words, for each theme, a stance would be (theme X positive; theme X negative; and theme X neutral). This is in keeping with earlier work on stance as a product of topic and sentiment (ALDayel and Magdy, 2021; Naskar et al., 2016).
3.3.4. Qualitative Analysis of Clusters
For each Louvain cluster, we sampled 10 comments from the most central users (by PageRank), double-coded them, and met to reach consensus on cluster character and central user roles. Initial categories (mentor, moderator, expert) were refined to include nuanced behaviors exhibited by those roles such as system critic and newcomer.
3.4. Classifier: Who are the cross-posters?
We framed cross-posting as a binary classification task with target cross-post $\in$ 0,1 representing whether the user posted to both r/maleinfertility and r/infertility.
We used the same feature set described in detail in §3.2.1. We framed cross-posting (posting to both r/maleinfertility and r/infertility) as a binary classification task. Using the same 95 features, we trained and evaluated Logistic Regression, Random Forest, Extra Trees, Gradient Boosting, HistGradientBoosting, AdaBoost, GaussianNB, KNN, LinearSVC, and XGBoost/LightGBM with 5-fold StratifiedKFold. The best model achieved accuracy = 0.846, precision = 0.793, recall = 0.307, F1 = 0.442, and ROC–AUC = 0.835, indicating strong fit (Lemeshow and Hosmer Jr, 1982; Fawcett, 2006; Swets, 2014). We applied SHAP to the final Gradient Boosting model to identify the most predictive features (Lundberg and Lee, 2017). Each SHAP value indicates a feature’s additive effect on the log-odds of cross-posting, with positive values increasing and negative values decreasing the likelihood of cross-posting (Lundberg and Lee, 2017).
3.5. Positionality and Ethical Stance
We are a team of researchers committed to supporting individuals navigating trauma, uncertainty, and life transitions. The first author, a cisgender man, studies how social technologies empower vulnerable communities while posing risks to privacy, autonomy, and identity. The second author, a cisgender South Asian woman, is a graduate student focused on reproductive and public health. The third author, a cisgender East Asian woman, examines how digital technologies can empower users. The last author, a nonparent white Latinx cisgender woman, specializes in transforming large, complex datasets into actionable insights for nonprofits, government, and industry, emphasizing the human stories behind the data.
Aware that online users often do not expect to be quoted directly (Fiesler and Proferes, 2018; Fiesler et al., 2024), we paraphrased community stances and edited supporting quotes to protect anonymity (Bruckman, 2002).
4. Findings
Our findings unfold across three layers of analysis. In §4.1, we present a thematic analysis showing how medical advice, emotional support, and moderation position the subreddit as a diagnostic hub, therapeutic commons, and digital institution, and how these practices influence future engagement. In §4.2, we examine community structure and roles along with how moderation and partner participation shape collective identity. Finally, §4.3 explores cross-posting, revealing how users who bridge r/maleinfertility and r/infertility serve as navigators and mentors, aligning norms across spaces and reinforcing male-centered support.
4.1. Diagnostic Utility Over Emotional Support: Themes Driving Future Engagement
Our analysis of 115 discussion topics, grouped under three overarching themes—Medical Advice, Emotional Support, and Moderation—illustrates how the infertility community on Reddit blends biomedical knowledge, affective support, and digital governance to sustain its members. Each theme encompasses multiple sub-themes that together reflect the community’s hybrid role as an informal medical forum, a therapeutic space, and a structured support system. Medical Advice (63.3%) centers on interpreting diagnostic tests, navigating treatment pathways, confronting advanced biological challenges, and addressing systemic barriers such as insurance. Emotional Support (7.4%) encompasses grief, coping rituals, comfort posting, and strategies for fostering resilience and social connection. Moderation (29.2%) highlights on-boarding rituals, navigation aids, rule enforcement, and the wiki/FAQ infrastructure that structures participation. Together, these themes show how the community functions simultaneously as an informal diagnostic hub, a therapeutic commons, and a regulated digital institution. A more detailed breakdown of the themes and topics can be seen in Appendix §A.2.
4.1.1. Medical Advice: Navigating Diagnostics, Treatments, and Systemic Barriers
The most prominent theme centers on medical advice and information exchange, where users collectively interpret results, evaluate treatment options, and contextualize both biological and systemic challenges.
Medical Report Analysis
A large cluster of discussions revolves around decoding semen analyses and other fertility test reports. Posters often admit confusion about terms like motility, morphology, or DNA fragmentation, with some worrying that a flagged number meant they were “doomed.” Peers step in to explain contested measures, reframing results as part of a broader diagnostic picture. For example, users emphasized that “percentages can mislead if counts are high,” reminding others that even a 0% morphology result did not rule out conception.
Female Infertility Generalized Guidance
Parallel threads focus on ovulation induction, endometrial lining preparation, or embryo transfer logistics. Posters compare regimens of Clomid, Letrozole, or estrace, and narrate experiences of thin lining or cycle cancellation. Discussions often reveal systemic frustrations: one user described “feeling like fertility treatment is another full-time job” due to scheduling battles and unclear protocols.
Male Infertility and Biological Limits
Discussions of azoospermia, low motility, and donor sperm highlight how male infertility intersects with identity and grief. Some describe the devastation of failed surgical retrievals, while others detail the emotional weight of using donor gametes—one poster admitted being more anxious about “telling family than actually using donor sperm.”
Health and Well-being
Beyond treatment, members share lifestyle strategies: supplements such as CoQ10, vitamin E, and L-carnitine, alongside advice to “ditch the laptop on the lap” or “switch to cooler underwear.” While highly anecdotal, these exchanges position men and women as active agents in shaping outcomes.
Advice on Insurance, Doctors, and Medication Expense
Another critical sub-theme is financial navigation. Posters swap tips on discount programs (e.g., Compassionate Care), employment benefits (e.g., Starbucks part-time IVF coverage), and shared-risk refund schemes. Others compare international systems, noting for example that “in the Netherlands, three cycles are covered with a deductible” whereas U.S. cycles often cost over \$10,000 out-of-pocket. This highlights inequities where access to Assistive Reproductive Technology (ART) depends as much on geography and insurance literacy as biology. When semen is only viable for analysis within several hours of taking a sample, users debated the merits of various brands of at-home testing kits when access to clinics (along with the discomfort of providing in-person samples) was limited.
Taken together, these sub-themes reveal Reddit as a secondary diagnostic and logistical space: members crowdsource interpretations, trade experiential knowledge, and provide rapid responses that often fill gaps left by inaccessible or opaque medical care.
4.1.2. Emotional Support: Coping, Resilience, and Connection
Running parallel to biomedical exchange is a robust theme of emotional support, where the subreddit functions as a therapeutic scaffold.
Grief and Disappointment
Threads often feature raw narratives following miscarriage, failed embryo transfers, or prolonged infertility. Posters describe feeling “numb,” “burned out,” or like they were “going through the motions of injections while emotionally checked out.” Community replies emphasize validation—reminding one another that every loss, even a chemical pregnancy, “is real and deserves mourning.” It is important to note that men who decide to have children with their partners are also expectant parents, sharing the full spectrum of emotions—joys and sorrows—throughout the process just as much as women do. These comments highlight that men feel the pain of infertility and the loss of a biological child just as deeply as their female counterparts.
Comfort Posting
Alongside grief are lighter posts that serve as comfort rituals. Members trade injection playlists, joke about “pin the tail on the donkey with my thighs,” or celebrate small victories like mastering self-administered shots. These threads blend humor with encouragement, reframing difficult procedures as milestones of resilience. However, some posts suggested that the poster should share their challenges with their therapist in addition to a show of some support.
Connection, Hope, and Celebration
The community also carves out space for tentative joy. Ritualized language such as “cautious congrats” during early betas tempers hope with shared recognition of risk. Others celebrate birthdays or procedural wins with “you guys rock!” affirmations, offering reprieve from clinical stress.
Talking to Family and Friends
A recurring pain point is how to navigate conversations with relatives and peers. Posters vent about insensitive pregnancy announcements or family pressure, while peers strategize about boundary-setting and selective disclosure. Advice often centers on protecting one’s “sanity first,” even if it means declining baby showers.
4.1.3. Moderation: Structure, Orientation, and Safety
A third theme highlights the importance of moderation and community norms in shaping participation. While these practices are often praised for maintaining safety and coherence, they also produce barriers and frustrations, especially for newcomers.
Moderation Policy
Automated bots and moderators enforce strict posting guidelines, regularly reminding users that “standalone culture is saved for complex topics.” Basic questions, introductions, or updates are redirected into daily or weekly threads. Some users express gratitude for this structure, noting that the rules keep the subreddit “focused and high-signal,” avoiding the clutter of repetitive or low-effort posts. Others, however, describe the system as alienating when they first join. One poster remarked that their message was “not visible until a mod approved it,” leaving them feeling like they were “posting into a void” at a moment of emotional urgency.
Navigation
Navigation challenges are a recurrent frustration. Newcomers often post in the wrong place and are redirected with templated bot messages pointing to the FAQ, wiki, or Welcome Wednesday thread. For some, these messages serve as useful orientation tools that provide quick access to resources; others describe them as “robotic and impersonal,” especially when received after disclosing sensitive struggles. A common critique was that automated responses “felt more like bureaucracy than support,” creating disorientation during moments of crisis. Yet, once acclimated, many members acknowledge that the structured system of daily treatment chats and thematic threads helps them find relevant discussions and reduces the chaos of a large community. This navigation is not only done using bot, but at times members of the community act as navigators as well.
Onboarding and Private Messaging
Onboarding is a blend of humor, empathy, and procedural instruction. Automated greetings often reference the subreddit as “the best shitty corner of the internet,” a self-aware attempt to soften the rigidity of rule enforcement. Many users appreciate this levity, interpreting it as a sign that moderators understand the emotional intensity of infertility. At the same time, others note that such humor does not offset the sting of rejection when their posts are removed. In some cases, moderators also encourage members to continue sensitive or off-topic conversations via private messages, a practice that both protects thread coherence and provides space for more personal exchanges. While some valued this redirection as an invitation to connect more deeply, others felt it reinforced exclusion by pushing vulnerable disclosures out of public view.
Table 1. LIWC coefficients predicting future engagement: Positive (left) vs. negative (right) linguistic marker
Positive LIWC Negative LIWC
Feature Coef StdErr t p Feature Coef StdErr t p
Affiliation 0.779 0.122 6.388 *** Adverb -0.350 0.067 -5.166 ***
Auxverb 0.256 0.075 3.435 *** Article -0.377 0.059 -6.348 ***
Cause 0.163 0.039 4.200 *** Death -0.049 0.022 -2.195 **
Certain 0.162 0.038 4.237 *** Family -0.206 0.055 -3.743 ***
Compare 0.218 0.043 5.029 *** Female -0.110 0.054 -2.051 ***
Conj 0.207 0.079 2.641 ** Focuspast -0.129 0.048 -2.682 ***
Discrep 0.086 0.041 2.092 ** Home -0.048 0.024 -2.010 ***
Ingest 0.119 0.046 2.599 *** Male -0.261 0.093 -2.809 ***
Insight 0.320 0.065 4.954 *** Money -0.121 0.025 -4.836 ***
Motion 0.091 0.034 2.673 *** Negate -0.188 0.034 -5.548 ***
Netspeak 0.079 0.034 2.308 ** Space -0.196 0.062 -3.179 ***
Prep 0.289 0.107 2.708 *** Tentative -0.155 0.074 -2.106 ***
Risk 0.121 0.051 2.393 ** We -0.381 0.078 -4.907 ***
Sad 0.214 0.043 4.963 ***
Sexual 0.072 0.030 2.397 **
Social 0.301 0.150 2.009 **
Swear 0.084 0.029 2.933 ***
Time 0.116 0.045 2.578 ***
Work 0.131 0.038 3.433 ***
4.1.4. Predictors of More Future Engagement
The time-lagged regression analysis presented in Tables 1-3 offers insights into how language use (sentiment), themes discussed by the user, and user engagement behavior on this subreddit community predict higher engagement from month to month in the intfertiility support community. Specifically, linguistic markers drawn from LIWC (Table 1) show that future engagement and language showing affiliation with others in the community, insight-oriented words, and cognitive processing terms are all positively associated with an increased engagement in the community. Conversely, language associated with detachment or externalization—such as references to money, negation, or fatalistic expressions—is negatively associated with continued participation.
Complementing the linguistic markers, thematic predictors (Table 2) show that users who contribute informational and strategic content—such as posts giving advice on insurance, Doctors, and Medication Expense and interpreting Medical Reports (see §4.1.1) as well Health and well-being posts are more likely to remain active in the community over time. In contrast, those whose contributions focus heavily on emotional processing (explained in §4.1.2) and include Comfort Posting or expressions of Connection, Hope, and Celebration, Grief and Disappointment or Talking to Family and Friends), are less likely to re-engage in the future. This may indicate that some users seek temporary emotional relief rather than ongoing participation. However, it might also indicate that social support might be continued on different channels. In our qualitative analysis, we noticed that a number of conversations where users are “venting” end in direct messages (DM) to one or more members of the community. They may also move these discussions to virtual meetings (e.g., Zoom meetings especially in the early days of the pandemic) or in-person meetings. While not focused on emotional processing, the Female Infertility Generalized Guidance also decreses the chance of future engagement in the community – probably because this community is more focused on Male Infertility Guidance.
Behavioral metrics from Table 3 reinforce this pattern: longer, more detailed posts and higher post scores (indicating community validation) are strong predictors of future activity, whereas highly positive sentiment or lack of replies correlates with disengagement. However, one surprising find is that the possibility of receiving a response has a negative relationship with the user’s continued engagement in the community. This may be related to moving the conversation to other spaces (as discussed above), or to the content of the responses which might demotivate users from future engagement. Finally, negative VADER sentiment is associated with increased future engagement, suggesting that users expressing more negative emotions are more likely to continue participating in the community. This aligns with the emotionally taxing nature of infertility—particularly for men, who often face unique challenges such as stigma, identity threat, limited offline support, and lower interest in male infertility in the scientific community. In this context, negative sentiment is not only expected but may reflect a critical need for support, validation, and shared understanding within a safe, anonymous space.
Taken together, the findings suggest that sustained engagement is driven not simply by emotional need, but by a sense of connection, informational utility, and perceived recognition. Community designers and moderators may foster deeper retention by encouraging bidirectional exchange, prioritizing actionable discussions, and ensuring contributors feel acknowledged and valued.
Table 2. Thematic predictors of future engagement: Positive (left) vs. negative (right) topic labels
Positive Theme Negative Theme
Feature Coef StdErr t p Feature Coef StdErr t p
Advice on Insurance, Doctors, and Medication Expense 0.051 0.024 2.136 *** Comfort Posting -0.040 0.016 -2.436 ***
Health and Well-being 0.061 0.024 2.591 ** Connection, Hope, and Celebration -0.065 0.030 -2.154 **
Medical Report Analysis 0.075 0.026 2.873 *** Female Infertility Generalized Guidance -0.054 0.024 -2.283 **
Grief and Disappointment -0.042 0.018 -2.380 **
Talking to Family and Friends -0.047 0.014 -3.309 ***
Table 3. Control variables predicting future engagement: Sentiment, score, comment characteristics, and response probability
Feature Coef StdErr t p
VADER Sentiment -0.108 0.0276 -3.904 ***
Score 0.072 0.020 3.685 ***
Comment Length 0.654 0.198 3.3 ***
Posibility of Response -0.103 0.041 -2.541 **
4.2. Community Structure and User Role Dynamics
After removing clusters with less than 10 users, the Louvain community detection algorithm provided us with eleven micro-communities – heretofore clusters. As shown in Table 4, all detected clusters exhibit conductance values below 0.5, suggesting that they are structurally cohesive and somewhat insulated from the broader network—a pattern that could indicate potential echo chamber behavior from a purely structural perspective. However, when we apply the second condition for echo chambers proposed by Morini et al. (Morini et al., 2021)—purity greater than 0.7—we find that none of the clusters meet this threshold, with purity values instead remaining low (ranging from 0.07 to 0.16). This indicates that while clusters maintain internal structural coherence, reinforced by moderate to substantial reciprocity, they do not exhibit the high topical or stance homogeneity necessary to be classified as echo chambers. Indeed, the relatively low purity scores, coupled with consistently high topic entropy values (3.9–5.2), suggest that each cluster engages with a wide diversity of themes and stances, a finding elaborated in Appendix B that shows the distribution of stances (defined in §3.3.3) across different clusters.
Indeed, we struggled to find thematic signatures for different clusters throughout our qualitative analysis of the clusters. With the exception of smaller clusters like cluster 2 made of 33 users, which focused on at-home sperm banking, most clusters showed a similar spread of varied topics related to male infertility. Even in that cluster, the discussion diverged to multiple different topics. However, similarities and differences emerged between different roles, and the ways in which they were enacted in different clusters. We discuss this in detail in §4.2.1.
Finally, as shown in Figure 1, most clusters experienced notable growth during the onset of the COVID-19 lockdowns in 2020, suggesting that broader social disruptions served as catalysts for increased participation in male infertility discussions.
Figure 1. Cluster membership growth over time. Most clusters show steady participation prior to 2020, followed by sharp increases during the COVID-19 lockdown period. This pattern suggests that broader social disruptions acted as catalysts for intensified engagement in male infertility discussions, contributing to the structural consolidation of clusters.
Table 4. Network cluster characteristics: Structural cohesion, reciprocity, and topical diversity metrics
Cluster Size Conductance Reciprocity Purity Topic
Entropy
11 2,614 0.313 0.279 0.13 4.907
5 1,852 0.270 0.328 0.08 4.829
4 1,594 0.312 0.312 0.07 5.232
15 1,013 0.365 0.334 0.09 4.667
6 953 0.378 0.347 0.07 4.926
13 877 0.397 0.345 0.07 4.775
7 800 0.378 0.346 0.09 4.718
27 59 0.326 0.375 0.14 4.522
2 33 0.364 0.255 0.10 4.568
23 26 0.420 0.390 0.12 3.918
18 15 0.222 0.346 0.16 3.899
Table 5. Distribution of user roles across network clusters
Cluster Roles
11 Mentor; Experts; Moderators; Systems Critic; Hope Broker
5 Mentor; Experts; Systems Critic; Hope Broker
4 Mentor; Hope Broker; Experts; Moderators
15 Mentor; Experts; Systems Critic; Hope Broker
6 Mentor; Experts; Systems Critic; Hope Broker; Newcomer
13 Mentor; Experts; Systems Critic; Hope Broker
7 Mentor; Experts; Systems Critic; Hope Broker; Newcomer
27 Mentor; Experts; Hope Broker; Newcomer
2 Mentor; Experts; Systems Critic; Hope Broker
23 Mentor; Experts; Systems Critic; Hope Broker; Newcomer
18 Mentor; Experts; Systems Critic; Hope Broker
4.2.1. Cross-Cluster Roles and Their Variation
Certain roles recur across clusters but manifest differently depending on the context. These roles are overlapping – often one poster is wearing “many hats” frequently within the context of a given thread as they balance providing medical information with socializing markers of commiseration. Users in these roles adopt certain commenting behaviors depending upon which role they are undertaking. Table 5 shows the breakdown of roles in different clusters.
Mentors, who represent the majority of the dominant voices within our clusters, are those participants who are ”shepherding” new and old members alike. Much of this guidance comes in the form of emotional support and advice based on their own lived experience. In Cluster 7, one mentor described in detail expected recovering sperm count thresholds after hormonal therapy, while in Cluster 6 another narrated their journey with acceptance that they would not have a child after multiple failed surgeries. In both examples, these Mentors modeled to other participants how they faced similar struggles with fertility: the former through detailed experiential medical information about recovery, and the latter through closure after failure. Mentors acting as hope brokers are nearly universal, sustaining morale across contexts: either by framing setbacks in terms of relatable shared experiences or by providing medical guidance to overcome obstacles. In Cluster 18, one partner emphasized improvements after months of supplements, while in Cluster 23 another highlighted dramatic gains following a vitamin regimen. Such accounts sustain optimism even when treatment outcomes remain uncertain.
Experts in different clusters offered detailed instructions on hormone regimens, supplements, and surgical timing. In Cluster 5, a user explained how Clomid influences hormonal pathways. In Cluster 11, another laid out strategies for recovering after testosterone therapy; and in Cluster 15, individuals logged precise hormone dosages and outcomes. Across clusters, experts provide structured, technical guidance using sometimes mechanistic or even clinical language, sprinkled occasionally with their anecdotal observations. Having extensive knowledge of an often opaque and byzantine landscape of navigating infertility health challenges, these users highlight the challenges all these users are facing. As such, comments that reflect their frustration with these health systems punctuate nearly every cluster. This systems critical behavior is often reflected by both the Expert and Mentor roles, as their first hand experience of the structural bias, financial burden, and professional blind spots makes their insights carry more weight. For instance, one user argued that fertility clinics often marginalize men, another criticized doctors for pressuring patients into surgery, and others warned about conflicts of interest in treatment recommendations. System critical comments thus shape the community’s positioning against the medical establishment, making visible the broader structures that frame individual infertility struggles. These communities exist largely because of the hardships their members have faced accessing and affording care as well as overcoming structural and societal biases regarding male infertility; it is no surprise that an undercurrent of systems criticism is so clearly visible.
4.2.2. Moderation by Users Across Clusters
Human moderators and everyday users enforce and interpret group norms via often explicit forum guidelines for post behavior. Some users apply the rules strictly, telling others that their medical results do not qualify for posting. Others soften the enforcement, explaining why certain parameters (such as morphology alone) are not considered meaningful when posting their medical records and asking for insights. Still others frame moderation through community response, pointing out how upvotes and engagement reflect shared perceptions of value of certain posts – e.g., posts by system critical users critiquing the medical establishment through the lens of their experiences. Together, these practices show how moderation is enacted both through both automation and social interpretation.
In Cluster 4, we identified an instance of norm-setting through empathy, as users caution against flaunting “good” results that might trigger others in the community. We found multiple examples of moderation taking the form of scientific correction, with members pushing back against poor medical advice and urging others to seek proper specialists. However, we found that Cluster 23 had little formal moderation; extreme claims, such as abstinence-based fertility hacks were not checked. Cluster 23 is one of the smallest identified clusters with 26 total users. This indicates that smaller clusters might have different moderation styles than the rest of the community.
Moderation thus emerges as a form of community guidance, balancing rule enforcement with “navigation labor.” This labor, often invisible, ensures that newcomers find their footing and that discussions remain accessible, high-signal, and emotionally safe. Importantly, mentors sometimes overlap with moderators in this function, herding discussions in ways that preserve both coherence and empathy.
Figure 2 summarizes the different roles we identified in our analysis. It shows that mentors and moderators do not only enforce posting standards, but actively perform way-finding by directing users to appropriate threads, providing structured resources, and sometimes intervening with empathetic language to soften redirects. By doing so, they serve as way-finders, moving discussions to appropriate threads, orienting newcomers, and preventing newcomers from feeling “lost” when first engaging in the subreddit. This shepherding role is not purely administrative—it blends procedural guidance with emotional counseling, often drawing on lived experiences to normalize confusion. For example, mentors not only provide medical or experiential knowledge but also explicitly guide members toward appropriate threads, FAQs, or resources, ensuring that vulnerable disclosures are not left unanswered. This highlights the dual role of mentors and moderators as both emotional anchors and navigators of community norms.
4.2.3. The Role of Women and Significant Others Across Clusters
Women and partners are especially visible in clusters where emotional processing and distrust of the medical system dominate. In Cluster 4, women acted as voices of empathy and shared grief. For example, one partner expressed sorrow over a recent miscarriage and directed others toward supportive communities, simultaneously affirming shared pain and providing peer guidance.
In Cluster 18, wives frequently narrate their husbands’ test results and frustrations with medical encounters. One described her partner’s severe semen analysis results, while another, in her capacity as a system critic recounted a doctor dismissing a possible connection between varicoceles (a medical condition that causes pain and swelling in the testicles) and infertility.
In other clusters, partners take on the expert role by being “partner-scientists.” They compile supplement regimens and interpret semen analysis results on behalf of their husbands. For example, one partner described how her husband’s sperm count continued to rise as he adhered to lifestyle changes, while another detailed an exact supplement stack that reportedly improved results. These examples illustrate how women take on the roles of record-keepers, advocates, and interpreters of their partners’ medical journeys in different clusters.
Figure 2. Roles and interactions within male infertility support clusters. The diagram illustrates how newcomers, mentors, moderators, and experts overlap in functions such as way-finding, emotional counseling, policy enforcement, evidence sharing, and strategic medical guidance. These roles collectively shape community navigation, knowledge exchange, and emotional resilience.
4.3. Cross-Posters as Navigators and Mentors in the Male Infertility Community
The results of classifier predicting cross-posting users who post to both r/maleinfertility and the main gender-mixed r/infertility are visualized using SHAP are presented in Figure 3. The details of these predictors are contextualized and explained below.
Cross-posters on Reddit can be characterized by both their activity patterns and linguistic style. They are generally highly active users, contributing more, longer comments than non-cross-posting users, with shorter average response times, and longer, more positive responses, indicating acceptace by other members of the community.
Comments by cross-posters usually focus on navigation, where mentors and moderators (see §4.2.2) help users to navigate (see §4.1.3 the norms of r/maleinfertility. Some of the moderators rely on their knowledge of the normative differences between r/maleinfertility and other subreddits to guide users through its use. In addition to guiding people in community navigation, cross-posting users also guided others in better understanding their medical reports (see §4.1.1).
Linguistically, cross-posters shift away from a self-focused voice (“I”) toward a collective or external orientation, using more “we” pronouns and male references while being less involved in female references. They also rely more on narrative or explanatory constructions with auxiliary verbs. These linguistic shifts align with their typical content focus: offering guidance on navigating the community and helping others interpret medical reports, as discussed above.
Taken together, cross-posters emerge as prolific, engagement-driven users who actively contribute to community support. They not only seek validation by posting across spaces but also provide mentorship and direction to others. Their emphasis on navigation suggests an effort to shape community norms in ways that better accommodate men—potentially filling gaps left unaddressed in other infertility forums.
Figure 3. Feature importance for predicting cross-posting behavior. The plot shows both the relative contribution of each feature to the model’s output (SHAP value magnitude) and the direction of influence. Higher values of average reply sentiment, total user comments, and linguistic markers such as male and “we” pronouns increase the likelihood of cross-posting, while self-focused language (“I”), temporal references, and high response volume reduce it.
5. Discussion
Our findings portray r/maleinfertility as a hybrid sociotechnical space that simultaneously functions as an informal diagnostic hub, a therapeutic commons, and a governed institution. This configuration helps men navigate the compounded uncertainties of male infertility—biological, relational, and psychological—while also working around Traditional Masculinity Ideologies (TMI) that discourage help-seeking and emotional expression (Messerschmidt, 2019; Cleary, 2012; Oliffe et al., 2023; Tsan et al., 2011). In contrast to the isolation and “perverse passivity” reported in clinical contexts (Dolan et al., 2017), the subreddit affords anonymity and asynchronicity that legitimize disclosure, echoing prior HCI/CSCW work showing that online communities can scaffold identity work and care practices for stigmatized groups (Ammari et al., 2025; Randazzo and Ammari, 2025; Andalibi, 2020). Below, we deepen this account in relation to our research questions (§5.1) and then translate the results into design (§5.2) and policy recommendations (§5.3).
5.1. Contributions to Current Literature
We organize our contributions around three analytic dimensions. First, affiliation, insight, and actionable guidance drive sustained engagement through informal diagnostics, logistical coordination, and therapeutic scaffolding. Second, cohesive micro-communities with adaptive, overlapping roles show how mentorship and moderation extend peer-support and platform-governance accounts. Third, cross-posters normalize norms and transfer know-how across communities, complicating models of knowledge brokering and collective resilience. Together these findings advance theories of social exchange, trauma-informed design, and platform governance by foregrounding the hybrid infrastructures and relational dynamics that sustain participation in stigmatized health contexts.
5.1.1. Affiliation, insight, and actionable guidance keep people coming back
To answer RQ1, we combined thematic analysis (115 topics; §4.1) and time-lagged regression (§4.1.4). Three discourse arenas emerged—Medical Advice (63.3%), Emotional Support (7.4%), and Moderation/Orientation (29.2%)—forming a hybrid infrastructure for diagnostics, therapy, and governance. Engagement rose with actionable guidance but declined when users received more replies, countering earlier findings that community responses foster future participation (Arguello et al., 2006; Lampe and Johnston, 2005; Backstrom et al., 2008; Chung et al., 2017; Lampe et al., 2010; Hwang and Foote, 2021). We unpack these patterns below.
Informal diagnostics and logistics.
The dominance of Medical Advice shows how peers translate lab values, treatment options, and system navigation into lay expertise that supplements—or compensates for—opaque clinical interactions. This extends work on socially sustained self-tracking and collective sensemaking, where personal data becomes shared, emotionally manageable, and actionable narratives (Li et al., 2010; Knittel et al., 2021; Mamykina et al., 2015; Choe et al., 2017; Lu et al., 2021; Chung et al., 2017; Gui et al., 2017). Social Exchange Theory explains why such disclosures persist despite stigma: intrinsic rewards (self-worth, reciprocity) and informational gains outweigh perceived risks (Emerson, 1972; Wu et al., 2023; Yan et al., 2016; Soc, 2011).
Therapeutic scaffolding
Emotional Support. Emotional Support spans grief processing and “comfort posting,” aligning with research on digital belonging and men’s mental health, where social connection mitigates depression and suicidality (McLaren and Challis, 2009; Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022; Steptoe et al., 2013). Asynchronous communication can likewise foster well-being (Hull et al., 2020; Rahman and Nayan, 2022) and support resilience and identity repair in trauma-adjacent settings (Ammari et al., 2025; Randazzo and Ammari, 2025). Yet our regression analysis complicates the common assumption that more replies drive retention: response probability negatively predicts subsequent activity. Possible explanations include overly corrective or impersonal feedback and off-platform migration (e.g., DMs, external groups), revising participation models in stigmatized contexts (Andalibi, 2020; Ammari et al., 2019).
While prior work highlights direct peer responses as central to support (Hwang and Foote, 2021)—a key tenet of trauma-informed design (Chen et al., 2022; Scott et al., 2023)—users may instead sustain engagement by lurking (Randazzo and Ammari, 2023) or distributing participation across multiple communities (Hwang and Foote, 2021; Ammari et al., 2025). We outline implications for these dynamics in the future work section.
5.1.2. Cohesive micro-communities with adaptive and overlapping roles
Community detection revealed structurally cohesive clusters that fell well below topical/stance purity thresholds, indicating they are not echo chambers (Morini et al., 2021). Within these clusters, shared and distinct roles emerged, expressed differently across groups.
Moderation as Both Safeguard and Flashpoint
r/maleinfertility is tightly moderated, as evidenced by its About section on the homepage which directed partners (of infertile men) to a dedicated thread. Moderation—both automated and human—guides information flow by redirecting posts to wikis or daily threads and safeguards the psychosocial well-being of community members by setting rules about acceptable content. Yet these same controls can alienate newcomers, and it remains unclear which moderation practices dampen engagement. This reflects the contextual nature of harm online (Gillespie, 2018): even well-intentioned “good news” posts may be discouraged when negative sentiment sustains participation (§4.1.4). This dynamic parallels the “pain olympics” among parents of children with special needs, where positive updates can provoke negative responses from peers facing deeper challenges (Ammari and Schoenebeck, 2015).
Way-Finding and Navigation Labor
While peer support is a recognized cornerstone of trauma-informed design (Chen et al., 2022; Scott et al., 2023), our findings show that support extends beyond upvotes, replies, or algorithmic curation (LaViolette and Hogan, 2019; Randazzo and Ammari, 2023). Mentors and moderators play critical way-finding roles: contextualizing rules, directing newcomers to curated resources, and reframing emotional disclosures into actionable steps. Building on platform governance and care-work scholarship, we identify an under-theorized “navigation labor,” where these actors blend rule enforcement with empathetic guidance.
Intersecting Roles and Evolving Peer Support
Other roles that overlap with moderation—particularly mentorship—complicate conventional models of peer support. The type of peer who engages with newcomers and how these interactions evolve shape both emotional safety and knowledge transfer, suggesting a fluid, multi-layered ecology of support.
5.1.3. Bridge-builders who normalize norms and transfer know-how
When answering RQ3, we found cross-posters resemble collective-oriented brokers: they write generalizable, guidance-seeking content, lean on “we” pronouns, and often activate when engagement is limited locally—complementing prior observations of cross-community knowledge brokering and empathy building especially when managing trauma (Ammari et al., 2025; Randazzo and Ammari, 2025). This finding also complicates the study of norm creation at different community levels (i.e., The Reddit platform $\rightarrow$ several related subreddits $\rightarrow$ one subreddit $\rightarrow$ micro-communities within the subrreddit) (Chandrasekharan et al., 2018) as it shows how cross-posters can engage in norm-setting in different communities relying on their deep knowledge of community norms in both spaces.
5.2. Design Implications
Grounded in the above, we reframe design not as content-topic recommendation but as role-aware, journey-sensitive infrastructure that supports safe visibility, effective way-finding, and portable care.
5.2.1. Role-transportability: Recommending Users Based on their Assumed Roles
Earlier work highlighted narrative transportability—providing users access to relevant content—offering design approaches to help people discover new communities and connect around shared topical interests (Randazzo and Ammari, 2023, 2025; Ammari et al., 2025).
We extend this idea to role transportability. Building on role embeddings that model structural similarity in networks (e.g., identifying users who bridge subgraphs) (Ahmed et al., 2020, 2019), we incorporate temporal dynamics (Grayson and Greene, 2019) and behavioral signals from discourse and interaction patterns (Saxena and Reddy, 2022; Akar, 2025). This hybrid approach produces role-based embeddings enabling recommendation systems to match newcomers with experienced members who have played relevant roles over time. Consistent with user agency (a central tenet of trauma-informed design as presented in §2.3.2) and affirmative consent design principles, individuals must retain the option to opt out of such recommendations (Im et al., 2021). This recommendation system would move from suggesting people who talked about X theme to “people who act like mentors/experts/seekers/hope brokers. Rather than pushing users into a single cluster by topic, the system would situationally connect them to interactional neighborhoods that match their current role needs (e.g., a new comer is surfaced to active mentors who recently provided high-quality diagnostic scaffolding). This better reflects how users likely “become part of a cluster” through interaction rather than content alone.
5.2.2. Therapist-ready, user-controlled summaries of on-platform engagement
Our data suggest that some disclosures move to DMs or off-platform, with many replies encouraging therapy (§4.1.2). To support continuity of care and reduce the emotional burden of “retelling” (Bagwell-Gray et al., 2015), we propose an opt-in system that summarizes and anonymizes a user’s recent online engagements (e.g., threads joined, key concerns, decisions), with editable and redactable outputs. These summaries could be exported to Fast Healthcare Interoperability Resources or other therapist–client media, complementing mixed therapy modalities such as two-way text communication (Hull et al., 2020).
5.2.3. Visible role and expertise tags—with optional professional verification
Building on our role findings in §4.2.1 and prior work on badges, awards, and verified participation (Burtch et al., 2022; Trujillo, 2022; Bornfeld and Rafaeli, 2017) in §2.3.2, we recommend a multilayered tagging and verification system to enhance visibility and trust within the community.
Community-earned role tags—such as Mentor, Navigator, Systems-wise, or Report-Interpreter—would be driven by behavioral signals and community nominations. These tags should be revocable and scoped to specific contexts so that recognition reflects sustained, relevant engagement rather than one-time actions.
Complementing these role tags, members could voluntarily adopt topic or skill tags (e.g., insurance navigation, TESE recovery) to make their expertise legible to seekers. In parallel, professionally verified accounts for clinicians and counselors would strengthen credibility, following verification approaches such as license checks and scoped disclaimers proposed in (Ammari et al., 2025). Importantly, these accounts must include clear guardrails to prevent their presence from being misconstrued as a substitute for clinical care.
5.3. Policy Implications.
Male infertility communities on Reddit reveal critical gaps in formal care: fathers often receive inadequate medical guidance and lack access to psychological support, fueling the system critic role that challenges medical and policy shortcomings. By formally recognizing moderation as care work, platforms and public health agencies should resource moderator training for health stigma contexts (trauma-informed scripts, de-escalation, misinformation triage). This acknowledges the dual governance/therapeutic role highlighted in our findings. Secondly, policy makers and health advocates alike could go further to bring awareness to male infertility in public health discourse. Campaigns and clinical guidelines should explicitly address men’s barriers under TMI, normalizing help-seeking and couple-oriented care pathways (Messerschmidt, 2019; Dolan et al., 2017; Seidler et al., 2016). Additionally, by mandating that mental health care be considered part of fertility care much of the stigma associated with that care as well as access to those invaluable resources could be made more available to this community of patients. As we’ve demonstrated, embed screening and referral to online/offline support alongside treatment planning, leveraging evidence that connection and belonging mitigate risk in men’s mental health (McLaren and Challis, 2009; Bailey et al., 2018; Holt-Lunstad and Steptoe, 2022).
6. Limitations and Future Work
This study faces several limitations. First, while time-lagged regression revealed that receiving replies can sometimes reduce the likelihood of future engagement, we are not yet certain why this occurs. One possibility is that conversations migrate to private messages or off-platform spaces, but it is also plausible that the tone or content of responses discourages sustained participation. This ambiguity calls for future qualitative research—particularly interviews with community members and moderators—to better understand the dynamics of responses and their impact on engagement.
Second, moderation emerged as both contentious and necessary. Automated enforcement and human moderators play essential roles in ensuring high-quality discourse, yet they are also described as alienating, especially by newcomers. The dual nature of moderation—as governance and as care work—warrants deeper exploration. Future studies could directly engage with moderators across infertility communities to map their challenges, strategies, and needs.
Finally, our network analysis provides only a snapshot of cluster structures. Modeling how clusters evolve over time, including their topical diversity and survival rates of users within them, would allow for a richer understanding of community trajectories. Longitudinal modeling could reveal how echo chamber risks, mentorship roles, and cross-posting practices shift during different stages of the infertility journey. Such work would deepen our grasp of how digital infertility communities adapt to changing membership, external disruptions (e.g., COVID-19 lockdowns), and evolving cultural discourses around masculinity and reproduction.
7. Conclusion
Male infertility communities on Reddit operate as informal diagnostic hubs, therapeutic commons, and governed spaces that help men navigate stigma and uncertainty. Role differentiation—especially mentors and moderators as way-finders—enables collective sensemaking while balancing care and rule enforcement. Linguistic and topical signals tied to affiliation and actionable guidance predict continued engagement, whereas some replies may inadvertently deter public return. Clusters are cohesive yet not echo chambers. Designing for navigation labor, empathetic moderation, and cross-community bridges can make support more inclusive and sustainable.
8. Acknowledgments
9. Appendices
Appendix A Research Methods
A.1. Topic Modeling
A.1.1. Appendix: Hyperparameter Selection
The hyperparameter search focused on three key parameters:
  1. (1)
    UMAP Neighborhood Size: The Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes et al., 2018) is used to reduce the dimensionality of the document embeddings before clustering. We varied the n_neighbors parameter across {20, 30, 40, 50}, which determines the number of neighboring sample points considered when constructing the manifold approximation.
    • Smaller values (e.g., 20) result in a more local view, capturing fine-grained structure and enabling smaller, tightly packed clusters.
    • Larger values (e.g., 50) promote a more global view, favoring larger, more cohesive clusters.
    • Adjusting this parameter helps balance local topic specificity and global topic cohesion.
  2. (2)
    Minimum Topic Size: The min_topic_size parameter controls the smallest allowable topic size in the clustering process (Grootendorst, 2020). We explored values of {20, 30, 40, 50}:
    • Lower values (e.g., 20) allow for the formation of many smaller topics, potentially increasing topic granularity.
    • Higher values (e.g., 50) enforce the creation of larger, more general topics, ensuring stability but possibly reducing topic diversity.
    • Setting this parameter appropriately is crucial, particularly for large datasets where higher values (e.g., 100–500) may be necessary to avoid excessive topic fragmentation.
  3. (3)
    N-gram Range: The n_gram_range parameter was varied between (1,2) and (1,3), affecting how words are tokenized in the topic representation using CountVectorizer (Pedregosa et al., 2011).
    • (1,2) n-grams allow for single words and bi-grams, capturing common phrase structures such as “climate change.”
    • (1,3) n-grams further include tri-grams, enabling more nuanced topic representations such as “machine learning model.”
    • This adjustment helps capture multi-word expressions and improves the semantic coherence of extracted topics.
A.1.2. Appendix: Model Evaluation
Each of the 32 model configurations (combining different UMAP neighborhood sizes, minimum topic sizes, and n-gram ranges) was trained and evaluated using the u_mass coherence score, a measure of topic quality that relies on word co-occurrence probabilities (Newman et al., 2010). Coherence scoring is widely used in topic modeling to assess how semantically meaningful the generated topics are. The configuration yielding the highest coherence score was selected as the optimal set of parameters for final topic modeling.
Figure 4. Figure shows the spread of topic models along with associated topics and coherence scores. The top model with the highest coherence score is highlighted along with its parameters.
A.2. Topic Breakdown
A.2.1. Appendix: Medical Advice Analysis
The Medical Advice cluster (63.34% of all discussion topics) represents the largest and most active domain of infertility discourse. It encompasses detailed exchanges on diagnostics, treatment planning, systemic barriers, and adjunct strategies. The treemap (Figure 5) shows how topics are distributed across major sub-clusters:
1. Advice on Insurance, Doctors, and Medication Expense
  • Key topics: 41, 94, 23, 100, 74
  • Discussions foreground financial and logistical barriers to care. Users share tips for contesting insurance denials, finding affordable clinics or pharmacies, and coping with high out-of-pocket costs.
  • Topic 41 focuses on accessing and managing fertility medications, where patients compare discounts, shared-risk programs, and international pharmacies.
  • These discussions reveal how medical advice is inseparable from the economic realities of fertility care.
2. Adoption, Foster Care, and Family Planning Alternatives
  • Key topics: 14, 25, 98, 35
  • Users weigh non-biological pathways such as adoption, donor eggs, and foster care.
  • Topic 14 emphasizes the emotional, ethical, and identity dilemmas of using donor gametes, while Topic 25 covers the decision-making process between IUI and IVF.
  • These posts articulate grief over lost genetic continuity alongside hope for alternative family-building.
3. Medical Report Analysis
  • Key topics: 10, 0, 90, 1
  • This cluster centers on interpreting semen analyses, hormone panels, and embryo grading.
  • Topic 10 covers deep dives into semen analysis and male factor infertility, while Topic 0 focuses on navigating varicocele and testicular treatments.
  • Topic 90 addresses treatment comparisons (IUI vs. IVF). Collectively, these illustrate how Reddit users act as peer translators of medical data, filling interpretive gaps left by clinicians.
4. Fertility, Medical Treatment Information
  • Key topics: 4, 13, 24
  • Focused on protocol-level knowledge, this sub-cluster includes embryo transfer preparation, ovulation induction, and questions about endometrial thickness.
  • Posters blend clinical metrics with emotional uncertainty, often sharing experiences of failed cycles and marginal results.
5. Generalized Health and Wellbeing
  • Key topics: 75, 68, 33, 18, 21
  • Highlights lifestyle interventions and everyday health practices.
  • Discussions include supplements (CoQ10, L-carnitine, antioxidants), exercise, diet, and abstinence from alcohol/porn.
  • While some advice is biomedical, much is experiential and anecdotal, reflecting a culture of self-experimentation alongside medical treatment.
Summary
This treemap illustrates the dominance of Medical Advice in infertility discourse, emphasizing how community members function as lay interpreters, navigators, and strategists. Medical data (semen analyses, hormone results) are contextualized through peer translation, while treatments are weighed against cost, accessibility, and systemic barriers. The presence of adoption and lifestyle sub-clusters shows how members also grapple with broader existential and self-care dimensions.
Figure 5. Hierarchical treemap of Medical Advice discussions (63.34% of topics). Subthemes include advice on insurance, doctors, and medication expenses; adoption and family planning alternatives; interpretation of medical reports; fertility treatment protocols; and generalized health and wellbeing. This cluster shows how community members act as lay interpreters of biomedical knowledge while navigating systemic and financial barriers.
A.2.2. Appendix: Emotional Support Treemap Analysis
The Emotional Support cluster (7.44% of all discussion topics) illustrates how the infertility community functions as a therapeutic scaffold, offering empathy, coping strategies, and collective resilience. The treemap (Figure 6) shows how topics are organized into several major sub-clusters:
1. Grief and Disappointment
  • Key topics: 8, 53, 40, 45, 54
  • Threads capture raw expressions of grief following miscarriages, failed IVF or IUI cycles, and prolonged infertility.
  • Members describe feelings of being “numb” or “broken,” often framing infertility as ambiguous loss.
  • Community responses include validation and refrains of solidarity (e.g., “you are not alone”), creating a shared lexicon of care.
2. Connection, Hope, and Celebration
  • Key topics: 9, 11, 62, 56
  • Discussions focus on moments of optimism: positive pregnancy tests, supportive clinic staff, or milestones achieved during treatment.
  • Humor, cautious celebration, and expressions of hope help balance the heavier narratives of grief.
  • These threads highlight how small wins become rituals of resilience.
3. Comfort Posting
  • Key topics: 17, 113, 89
  • Emotional check-ins where users seek comfort and reassurance.
  • Posts often serve as informal therapy, allowing members to vent frustrations, share daily struggles, and receive encouragement.
  • Community validation reinforces a sense of belonging during vulnerable moments.
4. Talking to Family and Friends
  • Key topic: 111
  • Focused on relational dynamics, this cluster captures the difficulty of sharing infertility experiences with family, friends, or partners.
  • Strategies include boundary-setting around pregnancy announcements, selective disclosure, or managing insensitive remarks.
  • These discussions underscore how infertility reshapes social relationships and communication practices.
5. Externalized Support Networks
  • Key topics: 110, 76, 6
  • Beyond family and peers, members discuss support from pets, online networks, and community rituals.
  • Pets are frequently described as emotional anchors, providing comfort and companionship when treatment outcomes disappoint.
  • These forms of externalized support highlight creative coping in the face of medical and social uncertainty.
Summary
The Emotional Support treemap illustrates the psychosocial dimension of infertility communities. Members engage in collective grieving, celebrate small victories, and co-create rituals of resilience. Through comfort posting and discussions of relational strain, the community offers not only emotional validation but also practical coping strategies for navigating stigma and isolation.
Figure 6. Hierarchical treemap of Emotional Support discussions (7.44% of topics). Subthemes include grief and disappointment, connection and celebration, comfort posting, talking to family and friends, and externalized support networks (e.g., pets, online ties). This cluster highlights the community’s role as a therapeutic space for collective grieving, resilience, and coping.
A.2.3. Appendix: Moderation Treemap Analysis
The Moderation cluster (29.22% of all discussion topics) highlights the digital governance and infrastructural practices that sustain community order, safety, and knowledge accessibility. The treemap (Figure 7) shows how moderation-related discussions are organized into several sub-clusters:
1. Navigation and Rule Guidance
  • Key topics: 29, 72, 52, 37, 19
  • These discussions focus on helping users find their way through the community’s rules, FAQs, and posting guidelines.
  • Members are often redirected to wiki pages, daily threads, or pinned posts when their questions or updates overlap with existing resources.
  • Navigation ensures high signal-to-noise ratio and maintains a structured flow of information, though it can sometimes frustrate newcomers.
2. Test Report Analysis and Posting Standards
  • Key topics: 20, 36
  • Moderators enforce rules around posting diagnostic data, such as semen analysis results or medical reports.
  • Bots and community guides provide templates for interpreting results, helping members avoid panic over borderline values.
  • These practices reflect trauma-informed moderation, balancing empathy with accuracy and order.
3. Moderation Policy and Enforcement
  • Key topics: 70, 39, 93, 78, 104
  • A cluster dedicated to enforcing participation guidelines and managing inappropriate posts.
  • Discussions cover redirecting updates to designated threads, discouraging repetitive questions, and maintaining tone.
  • This ensures respectful, high-quality discourse while preserving emotional safety for participants.
4. Onboarding and Community Orientation
  • Key topics: 83, 86, 3
  • New members are welcomed through automated but empathetic messages that introduce both rules and community culture.
  • Orientation blends practical guidance with humor, signaling that the space is simultaneously regulated and supportive.
  • This fosters early belonging while reinforcing boundaries and expectations.
Summary
The Moderation treemap illustrates the invisible labor of governance that underpins the infertility community. Through navigation guidance, posting standards, policy enforcement, and onboarding rituals, moderators and automated systems co-create a structured environment. This infrastructure not only protects informational quality but also enhances psychosocial safety, enabling the community to function as a reliable health and support resource.
Figure 7. Hierarchical treemap of Moderation discussions (29.22% of topics). Subthemes include navigation and rule guidance, test report posting standards, policy enforcement, and onboarding rituals. This cluster illustrates the community’s governance infrastructure, showing how moderators and bots maintain order, safeguard emotional safety, and structure knowledge exchange.
Appendix B Cluster Community Analysis
Appendix B presents the detailed stance-coding scheme referenced in §3.3.3. Tables 6-8 show each cluster defined in §4.2 along with the stances defined as themes (defined in §4.1) crossed with VADER sentiment.
Table 6. Stance distribution across clusters: Top 10 weighted theme-sentiment combinations by cluster
Cluster Top 10 Stances (with their weights)
11 1. Navigation (+ve) – 15.2%
2. Adoption Foster Care Family Planning Alternatives (+ve) – 12.1%
3. Medical Report Analysis (+ve) – 10.6%
4. Advice on Insurance Doctors and Medication Expense (+ve) – 9.1%
5. Advice on Insurance Doctors and Medication Expense (-ve) – 5.3%
6. Adoption Foster Care Family Planning Alternatives (-ve) – 5.3%
7. Navigation (-ve) – 5.3%
8. Generalized Health and Wellbeing (+ve) – 3.8%
9. Grief Disappointment (neu) – 3.0%
10. Grief Disappointment (-ve) – 3.0%
5 1. Advice on Insurance Doctors and Medication Expense (+ve) – 13.7%
2. Medical Report Analysis (+ve) – 10.0%
3. Adoption Foster Care Family Planning Alternatives (+ve) – 9.7%
4. Navigation (+ve) – 9.4%
5. Advice on Insurance Doctors and Medication Expense (neu) – 7.5%
6. Advice on Insurance Doctors and Medication Expense (-ve) – 4.8%
7. Test Report Analysis (+ve) – 4.8%
8. Navigation (neu) – 4.4%
9. Generalized Health and Wellbeing (neu) – 4.1%
10. Adoption Foster Care Family Planning Alternatives (-ve) – 4.0%
18 1. Medical Report Analysis (+ve) – 20.0%
2. Advice on Insurance Doctors and Medication Expense (+ve) – 16.4%
3. Grief Disappointment (neu) – 7.3%
4. Adoption Foster Care Family Planning Alternatives (neu) – 5.5%
5. Navigation (neu) – 5.5%
6. Advice on Insurance Doctors and Medication Expense (neu) – 3.6%
7. Grief Disappointment (+ve) – 3.6%
8. Medical Report Analysis (-ve) – 3.6%
9. Comfort Posting (-ve) – 3.6%
10. Health and Wellbeing (+ve) – 3.6%
Table 7. Stance distribution across clusters: Top 10 weighted theme-sentiment combinations by cluster (continued)
Cluster Top 10 Stances (with cluster weights)
4 1. Advice on Insurance Doctors and Medication Expense (+ve) – 12.9%
2. Adoption Foster Care Family Planning Alternatives (+ve) – 11.0%
3. Medical Report Analysis (+ve) – 10.8%
4. Navigation (+ve) – 9.6%
5. Advice on Insurance Doctors and Medication Expense (-ve) – 5.1%
6. Test Report Analysis (+ve) – 4.7%
7. Adoption Foster Care Family Planning Alternatives (-ve) – 4.5%
8. Moderation Policy (+ve) – 3.9%
9. Navigation (-ve) – 3.1%
10. Advice on Insurance Doctors and Medication Expense (neu) – 3.0%
15 1. Advice on Insurance Doctors and Medication Expense (+ve) – 14.5%
2. Medical Report Analysis (+ve) – 10.5%
3. Adoption Foster Care Family Planning Alternatives (+ve) – 9.9%
4. Navigation (+ve) – 9.2%
5. Advice on Insurance Doctors and Medication Expense (neu) – 6.1%
6. Test Report Analysis (+ve) – 5.8%
7. Advice on Insurance Doctors and Medication Expense (-ve) – 5.6%
8. Adoption Foster Care Family Planning Alternatives (-ve) – 3.6%
9. Adoption Foster Care Family Planning Alternatives (neu) – 3.4%
10. Navigation (-ve) – 3.1%
6 1. Adoption Foster Care Family Planning Alternatives (+ve) – 13.8%
2. Advice on Insurance Doctors and Medication Expense (+ve) – 11.0%
3. Medical Report Analysis (+ve) – 8.8%
4. Navigation (+ve) – 6.5%
5. Advice on Insurance Doctors and Medication Expense (-ve) – 5.9%
6. Test Report Analysis (+ve) – 5.7%
7. Adoption Foster Care Family Planning Alternatives (-ve) – 5.3%
8. Advice on Insurance Doctors and Medication Expense (neu) – 4.5%
9. Adoption Foster Care Family Planning Alternatives (neu) – 3.8%
10. Moderation Policy (+ve) – 2.9%
13 1. Advice on Insurance Doctors and Medication Expense (+ve) – 11.2%
2. Medical Report Analysis (+ve) – 10.5%
3. Adoption Foster Care Family Planning Alternatives (+ve) – 9.8%
4. Test Report Analysis (+ve) – 8.2%
5. Navigation (+ve) – 7.9%
6. Advice on Insurance Doctors and Medication Expense (neu) – 4.9%
7. Advice on Insurance Doctors and Medication Expense (-ve) – 4.5%
8. Adoption Foster Care Family Planning Alternatives (-ve) – 4.4%
9. Navigation (-ve) – 4.0%
10. Adoption Foster Care Family Planning Alternatives (neu) – 3.4%
Table 8. Stance distribution across clusters: Top 10 weighted theme-sentiment combinations by cluster (continued)
Cluster Top 10 Stances (with cluster weights)
7 1. Advice on Insurance Doctors and Medication Expense (+ve) – 14.9%
2. Medical Report Analysis (+ve) – 9.6%
3. Navigation (+ve) – 9.2%
4. Adoption Foster Care Family Planning Alternatives (+ve) – 9.1%
5. Test Report Analysis (+ve) – 6.5%
6. Advice on Insurance Doctors and Medication Expense (-ve) – 5.8%
7. Advice on Insurance Doctors and Medication Expense (neu) – 5.3%
8. Adoption Foster Care Family Planning Alternatives (neu) – 4.1%
9. Adoption Foster Care Family Planning Alternatives (-ve) – 4.1%
10. Navigation (-ve) – 3.8%
27 1. Advice on Insurance Doctors and Medication Expense (+ve) – 20.7%
2. Adoption Foster Care Family Planning Alternatives (+ve) – 12.6%
3. Medical Report Analysis (+ve) – 7.0%
4. Navigation (+ve) – 6.3%
5. Advice on Insurance Doctors and Medication Expense (-ve) – 6.0%
6. Moderation Policy (+ve) – 5.6%
7. Test Report Analysis (+ve) – 4.6%
8. Advice on Insurance Doctors and Medication Expense (neu) – 4.6%
9. Adoption Foster Care Family Planning Alternatives (neu) – 3.2%
10. Adoption Foster Care Family Planning Alternatives (-ve) – 3.2%
2 1. Navigation (+ve) – 15.2%
2. Adoption Foster Care Family Planning Alternatives (+ve) – 12.1%
3. Medical Report Analysis (+ve) – 10.6%
4. Advice on Insurance Doctors and Medication Expense (+ve) – 9.1%
5. Advice on Insurance Doctors and Medication Expense (-ve) – 5.3%
6. Adoption Foster Care Family Planning Alternatives (-ve) – 5.3%
7. Navigation (-ve) – 5.3%
8. Health and Wellbeing (+ve) – 3.8%
9. Grief Disappointment (neu) – 3.0%
10. Grief Disappointment (-ve) – 3.0%
23 1. Adoption Foster Care Family Planning Alternatives (+ve) – 12.3%
2. Test Report Analysis (+ve) – 12.3%
3. Medical Report Analysis (+ve) – 8.9%
4. Navigation (+ve) – 8.2%
5. Navigation (neu) – 8.2%
6. Advice on Insurance Doctors and Medication Expense (+ve) – 8.2%
7. Advice on Insurance Doctors and Medication Expense (neu) – 6.8%
8. Advice on Insurance Doctors and Medication Expense (-ve) – 4.8%
9. Test Report Analysis (-ve) – 3.4%
10. Navigation (-ve) – 3.4%